dont use online tools please
Use a double angle formula to find the exact value of cos ***
Graph the following equations: a. y = 4sin 2 (x + 1) b. y = 1 - cos x C. Y = 2 tanx + 1 d. y = 3csc x
Answer #1

Sols double angle formula Cos 11 A 12 Convert in to angle of - 165 Cos 165 Ces (gotas) = - sin (25) & -sin (45+30) I sinca +bgran, (a) y = 4 sin a (n + 1/ ) amplitude a = 4 Period I 2TL (b) ya in Cosne S(6) Y= 2 tann tl A (d) y a 3 cosecn -1 In - 7T -34 T LI IT